I posted this problem before and the answer was incorrect, so please can an expert with great knowledge help me this time.

(25) 4. Find the solution to the given system that satisfies the given initial c [1 0 -17 X = 0 2 0 X(t), X(0) = [1 0 1
Answer #1

X X(0) x lt) 2 2 0 Let becomes thon -bc (0) -2 21 21 22 - TovinLoploctranfarm in beth Side uotion the abovs Sx(s-o) = Xls -(sor, x(S = 2. (s-) (s-) Tovina inverse Loploce transform brih side , Sinht 2さCosht - From (3) oe S-1)) Substitutiog volue x) iN.B. sinh(t), cosh(t) are hyperbolic function .

If you have any doubt or need more clarification at any step please comment.